OpenStudy (anonymous):

What are the vertices of the hyperbola given by the equation 16x2 – y2 + 96x + 10y + 103 = 0? (-2, 5) and (-4, 5) (1, 5) and (-7, 5) (-3, 6) and (-3, 4) (-3, 9) and (-3, 1)

OpenStudy (anonymous):

Someone Please Help!!!!

OpenStudy (anonymous):

Try to complete the square 16x^2 + 96x + 144 = 16(x + 3)^2 y^2 + 10y + __ = (-y + 5)^2

OpenStudy (anonymous):

confused... I dont know how.

OpenStudy (anonymous):

on the wrong track, sorry

OpenStudy (anonymous):

can you help me? Do you know how to answer this?

OpenStudy (anonymous):

I think what you need is a form like \[\frac{(y-k)^2}{a^2} - \frac{(x-h)^2}{b^2} = 1\]

OpenStudy (anonymous):

That's why I was trying to complete the square (and did, for x). I ran into a little trouble with y.

OpenStudy (anonymous):

What did you get for x?

OpenStudy (anonymous):

16x^2 + 96x + 144 = 16(x + 3)^2

OpenStudy (anonymous):

Can you plug x back into the equation and find y?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Another possibility is to try the proposed solutions. If you plug in (-2,5) do you get 0?

OpenStudy (anonymous):

I will try all of them, and let you know what works!!!!

OpenStudy (anonymous):

Clearly (1, 5) and (-7, 5) doesn't work b/c the first term is definitely not a zero.

OpenStudy (anonymous):

It is the first one, Thank you!!!

OpenStudy (anonymous):

yw